# MultiMatch: Multihead Consistency Regularization Matching for Semi-Supervised Text Classification

### v2

@@ 5 Results
The comparison between MultiMatch and other SSL methods on the text classification datasets included in the USB benchmark is presented in Table 1. Our results show that MultiMatch consistently outperforms all the baselines, achieving top performance in 9 out of 10 setups and ranking first according to the Friedman test among 19 SSL algorithms. While the overall mean error also decreases by an absolute value of 0.28% compared to the second best approach, we notice that the improvement is higher when the number of labeled examples is lower. Specifically, the improvement over MarginMatch is on average 0.43% in the lower-resource setting, compared to 0.14% when more labeled data is provided.
By analyzing the baselines that provided inspiration for our approach, we observe that MarginMatch consistently outperforms all methods except MultiMatch, reinforcing the importance of leveraging historical evidence for pseudo-labels filtering. In contrast, Multihead Co-training and FreeMatch rank $12^{th}$ and $8^{th}$, respectively, indicating that the voting mechanism and the self-adaptive thresholds, while beneficial, are not among the most impactful techniques when used in isolation. However, when refined and incorporated into our unique three-fold pseudo-label weighting module (PLWM), their combined effect leads to substantial performance gains across all evaluation setups.
@@
Imbalanced settings.
MultiMatch outperforms all the baselines, ranking first in 7 out of 10 setups and outperforming the second-best approach, FreeMatch, by 3.26% averaged across all tasks. In particular, MultiMatch reaches state-of-the-art performance in all five setups with the imbalance factor 100, when the unlabeled distribution is similar to the labeled one, and surpasses FreeMatch by 3.48% when the unlabeled distribution is different. Comparing to MarginMatch, the main competitor in the balanced setup, MultiMatch surpasses it by an average of 8.1% in the imbalanced setting, with an impressive 11.74% improvement when the unlabeled distribution differs from the labeled one. A detailed technical discussion explaining how the design choices in MultiMatch contribute to its substantial performance gains is provided in Appendix B.
@@ 6 Ablation Study and Analysis
First, in order to assess the importance of useful & difficult examples, we remove the weight $w_{d}$, so they will be weighted similar to the useful & easy samples. This leads to a drop in performance on 9 out of 10 setups, with a mean error rate increase of 0.3%. Similarly, by completely removing these examples from the training data, on average we do not notice any substantial differences. Second, we remove the filtering based on the self-adaptive thresholds that rely on the current confidence of the model, $\mathbb{1}_{FreeMulti}$. This results in our PLWM using only the heads agreement and the APM score for filtering pseudo-labels. This leads to a significant drop in performance of 1.65%, highlighting the importance of current model’s beliefs in our PLWM. Third, we remove the APM-based filtering $\mathbb{1}_{Multi}$ from MultiMatch, obtaining a PLWM module that only uses the agreement samples where at least one head has high current confidence in the selected pseudo-label. In this case, the mean error rate increases by 1.21%, and the Friedman test places it the lowest among all the setups in the ablation study. Finally, removing the low limit for the class-wise thresholds $\gamma_{k,c}>0$ has the lowest impact of only 0.15% error rate, but it still provides additional proof that a stricter filtering mechanism is essential for reducing the noise in the pseudo-labels and preventing their degradation over time.
Figure 2: Mask rate (left) and impurity (right) on AG News with 200 labels.
Pseudo-Label Quality Analysis.
We compare MultiMatch with MarginMatch, Multihead Co-training, and FreeMatch in terms of pseudo-label quality using two metrics: mask rate and impurity. We show the results on the AG News task with 200 labels and its highly imbalanced version in Figures 2 and 3, respectively. The mask rate is defined as the proportion of pseudo-labeled examples that were excluded from training at epoch $t$, while impurity is the proportion of pseudo-labeled examples that participate in training at epoch $t$ but have an incorrect label. As expected, our PLWM module employs a stricter filtering mechanism, resulting in a higher mask rate than all other approaches. In contrast, relying only on heads agreement proves ineffective, as Multihead Co-training exhibits a mask rate of less than 1%. Meanwhile, our strict filtering, combined with a targeted approach for handling difficult examples, effectively minimizes impurity in both setups. In contrast, Multihead Co-training incorporates many erroneous pseudo-labels because of its weak filtering, while the self-adaptive thresholds of FreeMatch fail to prevent error accumulation in the imbalanced setup, causing impurity to rise as training progresses.
7 Conclusion
Limitations
While MultiMatch demonstrates strong performance across various setups, it has certain limitations. First, its reliance on multiple classification heads increases computational overhead compared to single-head SSL methods. In our case, the number of parameters increased by roughly 1% (from 110M to  [^1] M), but this may vary depending on the backbone. However, this overhead is negligible compared to traditional co-training approaches that train two distinct models. We extend the efficiency considerations in Appendix F. Second, applying MultiMatch to specialized applications may require domain-specific knowledge integration (e.g., customized augmentation strategies) and additional hyperparameter tuning. However, this limitation is shared by all SSL methods. In Appendix H, we take initial steps toward assessing the generalizability of MultiMatch by applying it in a multimodal, real-world experimental setting. The results highlight the particular robustness of MultiMatch across two distinct tasks, even without any hyperparameter tuning, in contrast to other baseline methods. Nonetheless, further analysis remains advisable for highly sensitive applications.
Appendix A Class-imbalanced Semi-supervised Learning
Appendix B MultiMatch for Class Imbalance
Appendix C Extended Background
with $\tau_{0}=\tilde{p}_{0}(c)=\frac{1}{C}$.
$\mathbb{1}_{Free}^{(t)}=\mathbb{1}(\max(q_{b})>\tau_{t}(\hat{q_{b}}))$ (10)
Appendix D Experimental Setup
Shared Hyperparameter Settings.
MultiMatch Hyperparameters.
The hyperparameters specific to MultiMatch are the number of heads $H$, which we set to 3 from the start, similar to Multihead Co-training, the weight $w_{d}$ used for the useful & difficult samples and the percentile $f$ used in computing our self-adaptive APM thresholds. For the APM percentile $f$ we experimented with values of 5% and 10%, and selected 5% as the optimal value. Setting $f$=10determined an increased mask rate, which excluded relevant samples from the unsupervised loss computation. For the weight $w_{d}$ of the useful & difficult examples, we experimented with values of 3 and 5, selecting 3 as the optimal value. The hyperparameters selection was done following experiments on the AG News dataset with 200 labels and choosing the value that resulted in a smaller validation error averaged over 3 runs. The method was chosen due to the faster training time on AG News, but it does not exclude that different values might be optimal for other tasks. However, our selected values provided consistent performance across all setups.
Computational Resources.
We conduct our experiments on a system with 8 Nvidia H100 GPUs with 80GB VRAM. Each experiment was run in a single GPU setting, the multiple GPUs being used to run multiple experiments simultaneously. Each training run took between 2 and 4 hours depending on the dataset ($\approx$2h for AG News, $\approx$3h for IMDB and Yahoo, and $\approx$4h for Amazon and Yelp). With an average of $\approx$3h per run, 3 runs per setup, 10 setups (5 datasets, 2 sizes each) and 3 models (MultiMatch, MarginMatch, Multihead Co-training), the total training time for the main results table is about 3x3x10x3 = 270 GPU hours. For the imbalanced setups we trained 10 models (5 models with and without ABC) for a total of 3x3x10x10 = 900 GPU hours). The ablation study required to train 5 additional models for a total of 3x3x10x5 = 450 GPU hours. Adding this up, the total cost was about 1,620 GPU hours, or 67.5 GPU days.
Appendix E Datasets
Dataset Label Type #Classes #Labeled #Unlabeled #Validation #Test
IMDB Movie Review Sentiment 2 10 / 50 11,500 1,000 12,500
AG News News Topic 4 10 / 50 25,000 2,500 1,900
Amazon Review Product Review Sentiment 5 50 / 200 50,000 5,000 13,000
Yahoo! Answer QA Topic 10 50 / 200 50,000 5,000 6,000
Yelp Review Restaurant Review Sentiment 5 50 / 200 50,000 5,000 10,000
Appendix F Efficiency Considerations
Appendix G Results
Appendix H Multimodal Setup
To evaluate the generalizability of our approach in real-world multimodal scenarios, we apply MultiMatch to the multimodal version of the CrisisMMD dataset Ofli et al. (2020) and compare its performance against several representative baselines. We adopt the experimental setup proposed by Sirbu et al. (2022), which includes access to unlabeled data suitable for semi-supervised learning. Specifically, we use MMBT Kiela et al. (2019) as the supervised backbone and follow the same hyperparameter configuration as in their work. Importantly, we apply MultiMatch with the same hyperparameters used in the NLP experiments, without any task-specific tuning.
We selected CrisisMMD for several reasons. First, it provides real-world data from the crisis response domain. Second, the dataset is multimodal – combining text and images – unlike the previously used NLP datasets. Third, it supports multiple classification tasks with distinct characteristics: for example, the Informative task features a nearly balanced binary classification setup, while the Humanitarian task involves five classes and is highly imbalanced. Additional details on the dataset can be found in Ofli et al. (2020).
As shown in Table 5, MultiMatch demonstrates strong generalization, consistently outperforming all other SSL baselines across both tasks. In contrast, other methods exhibit inconsistent performance. For instance, FreeMatch and Multihead Co-training perform well on the Informative task but struggle on the Humanitarian task. MarginMatch, which previously showed competitive results on NLP datasets, is outperformed by several alternatives in this multimodal setting. While a deeper investigation into domain-specific challenges lies outside the scope of this work, the results highlight the robustness of MultiMatch in multimodal, domain-specific, and highly imbalanced setups.
Although the exploration of additional real-world applications is beyond the scope of this work, the findings in this section highlight the strong potential of MultiMatch as a general-purpose SSL method. It performs reliably even in domain-specific and multimodal settings – without the need for task-specific adjustments or hyperparameter tuning – supporting its use as an off-the-shelf solution in diverse low-resource scenarios.
@@ Appendix H Multimodal Setup
Appendix I Code and License
All our code will be shared with an MIT license, similar to the USB framework Wang et al. (2022). We aim to integrate MultiMatch in USB as an open-source contribution.
@@ Appendix I Code and License
10.59 ±0.66 44.43 ±0.98 42.09 ±0.28 30.90 ±0.70 29.39 ±0.39 42.16 ±0.79 39.08 ±0.55 26.37 1.10 1
Table 6: Test error rates on IMDB, AG News, Amazon Review, Yahoo! Answer, and Yelp Review datasets using two setups with different sizes for the labeled set. The best result for each setup is highlighted in blue.
@@
Pseudo-Labeling Lee et al. (2013) 37.20 15.30 16
Mean Teacher Tarvainen and Valpola (2017) 33.13 14.10 14
$\Pi$-Model Rasmus et al. (2015) 51.13 16.60 17
VAT Miyato et al. (2018) 32.50 11.90 13
MixMatch Berthelot et al. (2019b) 36.31 14.30 15
ReMixMatch Berthelot et al. (2019a) 75.00 18.90 19
AdaMatch Berthelot et al. (2021) 27.59 5.70 3
UDA Xie et al. (2020a) 60.19 17.60 18
FixMatch Sohn et al. (2020) 29.83 8.70 10
FlexMatch Zhang et al. (2022) 28.23 8.20 8
Dash Xu et al. (2021) 30.33 10.30 11
CRMatch Fan et al. (2023) 27.69 6.40 5
CoMatch Li et al. (2021) 27.94 7.10 7
SimMatch Zheng et al. (2022) 27.94 6.60 6
FreeMatch Wang et al. (2023) 28.21 8.40 9
SoftMatch Chen et al. (2023) 27.42 5.70 3
Multihead Co-training Chen et al. (2021) 29.49 10.50 12
MarginMatch Sosea and Caragea (2023) 26.65 2.00 2
MultiMatch Ours 26.37 1.10 1

### v3

@@ 5 Results
The comparison between MultiMatch and other SSL methods on the five text classification datasets is presented in Table 1. Our results show that MultiMatch consistently outperforms all the baselines, achieving top performance in 8 out of 10 setups and ranking first according to the Friedman test among 21 SSL algorithms. While the overall mean error also decreases by an absolute value of 0.28% compared to the second best approach, we notice that the improvement is higher when the number of labeled examples is lower. Specifically, the improvement over MarginMatch is on average 0.43% in the lower-resource setting, compared to 0.14% when more labeled data is provided.
By analyzing the baselines that inspired our approach, we observe that MarginMatch consistently outperforms all methods except MultiMatch, reinforcing the importance of leveraging historical evidence for pseudo-labels filtering. In contrast, Multihead Co-training and FreeMatch have a lower rank, $14$ and $10$, respectively, indicating that the voting mechanism and the self-adaptive thresholds, while beneficial, are not among the most impactful techniques when used in isolation. However, when refined and incorporated into our unique three-fold pseudo-label weighting module (PLWM), their combined effect leads to substantial performance gains across all evaluation setups.
Imbalanced settings.
@@
MultiMatch outperforms all the baselines, ranking first in 7 out of 10 setups and outperforming the second-best approach, FreeMatch, by 3.26% averaged across all tasks. In particular, MultiMatch reaches state-of-the-art performance in all five setups with the imbalance factor 100, when the unlabeled distribution is similar to the labeled one, and surpasses FreeMatch by 3.48% when the unlabeled distribution is different. Comparing to MarginMatch, the main competitor in the balanced setup, MultiMatch surpasses it by an average of 8.1% in the imbalanced setting, with an impressive 11.74% improvement when the unlabeled distribution differs from the labeled one.
The superior performance of MultiMatch in highly imbalanced settings can be attributed to several CISSL techniques incorporated into its design. First, the use of class-wise thresholds enables the algorithm to account for the distinct learning progress of individual classes. For the same reason, FreeMatch shows a substantial improvement in imbalanced setups, ranking second before the ABC enhancement. Second, the use of percentile-based thresholds mitigates imbalance amplification, as shown by Guo and Li (2022). Finally, MultiMatch explicitly promotes lower pseudo-label impurity through several mechanisms. The use of multiple filters increases the difficulty for incorrect pseudo-labels to pass, while the subset employed for threshold computation results in generally higher thresholds compared to MarginMatch. This approach helps mitigate long-term error accumulation, a key challenge in imbalanced SSL. For a more detailed technical discussion, see Appendix B.
@@ 6 Ablation Study and Analysis
First, in order to assess the importance of useful & difficult examples, we remove the weight $w_{d}$, so they will be weighted similar to the useful & easy samples. This leads to a drop in performance on 9 out of 10 setups, with a mean error rate increase of 0.3%. Similarly, by completely removing these examples from the training data, on average we do not notice any substantial differences. Second, we remove the filtering based on the self-adaptive thresholds that rely on the current confidence of the model, $\mathbb{1}_{FreeMulti}$. This results in our PLWM using only the heads agreement and the APM score for filtering pseudo-labels. This leads to a significant drop in performance of 1.65%, highlighting the importance of current model’s beliefs in our PLWM. Third, we remove the APM-based filtering $\mathbb{1}_{Multi}$ from MultiMatch, obtaining a PLWM module that only uses the agreement samples where at least one head has high current confidence in the selected pseudo-label. In this case, the mean error rate increases by 1.21%, and the Friedman test places it the lowest among all the setups in the ablation study. Finally, removing the low limit for the class-wise thresholds ($\gamma_{min}$) has the lowest impact of only 0.15% error rate, but it still provides additional proof that a stricter filtering mechanism is essential for reducing the noise in the pseudo-labels and preventing their degradation over time.
Figure 2: Mask rate and impurity on AG News with 200 labels.
@@
Pseudo-Label Quality Analysis.
We compare MultiMatch with MarginMatch, Multihead Co-training, and FreeMatch in terms of pseudo-label quality using two metrics: mask rate and impurity. We show the results on the AG News task with 200 labels in Figure 2 (similar results are shown in Figure 3 in Appendix B for the imbalanced setting). The mask rate is defined as the proportion of pseudo-labeled examples that were excluded from training at epoch $t$, while impurity is the proportion of pseudo-labeled examples that participate in training at epoch $t$ but have an incorrect label. As expected, our PLWM module employs a stricter filtering mechanism, resulting in a higher mask rate than all other approaches. In contrast, relying only on heads agreement proves ineffective, as Multihead Co-training exhibits a mask rate of less than 1%. Meanwhile, our strict filtering, combined with a targeted approach for handling difficult examples, effectively minimizes impurity in both setups. In contrast, Multihead Co-training incorporates many erroneous pseudo-labels because of its weak filtering, while the self-adaptive thresholds of FreeMatch fail to prevent error accumulation in the imbalanced setup, causing impurity to rise.
7 Performance in Multimodal Settings
We evaluate the generalizability of MultiMatch in real-world multimodal settings by applying it to the multimodal version of the CrisisMMD dataset Ofli et al. (2020) and compare its performance against several representative baselines. We adopt the experimental setup from Sirbu et al. (2022), which includes access to unlabeled data suitable for SSL. Specifically, we use MMBT Kiela et al. (2019) as the supervised backbone and follow the same hyperparameter configuration as in their work. Importantly, we apply MultiMatch with the same hyperparameters used in the NLP experiments, without any task-specific tuning.
We selected CrisisMMD for several reasons. First, it provides real-world data from the crisis response domain. Second, the dataset is multimodal – text and images – unlike the previously used NLP datasets. Third, it supports multiple classification tasks with distinct characteristics: for example, the Informative task features a nearly balanced binary classification setup, while the Humanitarian task involves five classes and is highly imbalanced. Additional data details can be found in Ofli et al. (2020).
As shown in Table 4, MultiMatch demonstrates strong generalization, consistently outperforming all other SSL baselines across both tasks. In contrast, other methods exhibit inconsistent performance. For instance, FreeMatch and Multihead Co-training perform well on the Informative task but struggle on the Humanitarian task. MarginMatch, which previously showed competitive results on NLP datasets, is outperformed by several alternatives in this multimodal setting. While we do not delve into a deeper investigation of domain-specific challenges, the results highlight the robustness of MultiMatch in multimodal, domain-specific, and highly imbalanced setups.
8 Conclusion
Acknowledgements
This research is supported by the project “Romanian Hub for Artificial Intelligence - HRIA”, Smart Growth, Digitization and Financial Instruments Program, 2021-2027, MySMIS no. 334906; the NSF IIS award 2107518; and a UIC Discovery Partners Institute (DPI) award. Any opinions, findings, and conclusions expressed here are those of the authors and do not necessarily reflect the views of NSF or DPI. We also thank our anonymous reviewers for their insightful feedback and suggestions.
Limitations
While MultiMatch demonstrates strong performance across various setups, it has certain limitations. First, its reliance on multiple classification heads increases computational overhead compared to single-head SSL methods. In our case, the number of parameters increased by roughly 1% (from 110M to  [^1] M), but this may vary depending on the backbone. However, this overhead is negligible compared to traditional co-training approaches that train two distinct models. We extend the efficiency considerations in Appendix F. Second, applying MultiMatch to specialized applications may require domain-specific knowledge integration (e.g., customized augmentation strategies) and additional hyperparameter tuning. However, this limitation is shared by all SSL methods. In Section 7, we take initial steps toward assessing the generalizability of MultiMatch by applying it in a multimodal, real-world experimental setting. The results highlight the particular robustness of MultiMatch across two distinct tasks, even without any hyperparameter tuning, in contrast to other baseline methods. Nonetheless, further analysis remains advisable for highly sensitive applications.
Appendix A Class-imbalanced Semi-supervised Learning
Appendix B MultiMatch for Class Imbalance
Appendix C Extended Background
with $\tau_{0}=\tilde{p}_{0}(c)=\frac{1}{C}$.
$)\mathbb{1}_{Free}^{(t)}=\mathbb{1}(\max(q_{b})>\tau_{t}(\hat{q_{b}}$)) (10)
Appendix D Experimental Setup
Shared Hyperparameter Settings.
MultiMatch Hyperparameters.
The hyperparameters specific to MultiMatch are the number of heads $H$, which we set to 3 from the start, similar to Multihead Co-training, the weight $w_{d}$ used for the useful & difficult samples, the percentile $f$ used in computing our self-adaptive APM thresholds, and their lower bound $\gamma_{min}$. For the APM percentile $f$ we experimented with values of 5% and 10%, and selected 5% as the optimal value. Setting ff=10% determined an increased mask rate, which excluded relevant samples from the unsupervised loss computation. For the lower bound $\gamma_{min}$, we experimented with the values 0 and -$\infty$ (no lower bound), and found that -$\infty$ works better in highly imbalanced setups as it allows for more pseudo-labels from the underrepresented classes. For the weight $w_{d}$ of the useful & difficult examples, we experimented with values of 3 and 5, selecting 3 as the optimal value. The hyperparameters selection was done following experiments on the AG News dataset with 200 labels and choosing the value that resulted in a smaller validation error averaged over 3 runs. The method was chosen due to the faster training time on AG News, but it does not exclude that different values might be optimal for other tasks. However, our selected values provided consistent performance across all setups.
Dataset Label Type #Classes #Labeled #Unlabeled #Validation #Test
IMDB Movie Review Sentiment 2 10 / 50 11,500 1,000 12,500
AG News News Topic 4 10 / 50 25,000 2,500 01,900
Amazon Review Product Review Sentiment 5 50 / 200 50,000 5,000 13,000
Yahoo! Answer QA Topic 10 50 / 200 50,000 5,000 06,000
Yelp Review Restaurant Review Sentiment 5 50 / 200 50,000 5,000 10,000
Computational Resources.
We conduct our experiments on a system with 8 Nvidia H100 GPUs with 80GB VRAM. Each experiment was run in a single GPU setting, the multiple GPUs being used to run multiple experiments simultaneously. Each training run took between 2 and 4 hours depending on the dataset ($\approx$2h for AG News, $\approx$3h for IMDB and Yahoo, and $\approx$4h for Amazon and Yelp). With an average of $\approx$3h per run, 3 runs per setup, 10 setups (5 datasets, 2 sizes each) and 5 models (MultiMatch, MarginMatch, Multihead Co-training, SequenceMatch, CGMatch), the total training time for the main results table is about 3x3x10x5 = 450 GPU hours. For the imbalanced setups we trained 10 models (5 models with and without ABC) for a total of 3x3x10x10 = 900 GPU hours). The ablation study required to train 5 additional models for a total of 3x3x10x5 = 450 GPU hours. Adding this up, the total cost was about 1,800 GPU hours, or 75 GPU days.
Appendix E Datasets
Appendix F Efficiency Considerations
Appendix G Results
@@
10.59 ±0.66 44.43 ±0.98 42.09 ±0.28 30.90 ±0.70 29.39 ±0.39 42.16 ±0.79 39.08 ±0.55 26.37 01.2 01
Table 6: Test error rates on IMDB, AG News, Amazon Review, Yahoo! Answer, and Yelp Review datasets using two setups with different sizes for the labeled set. The best result for each setup is highlighted in blue. The upper section of the table contains results reported by the USB benchmark, while the bottom section contains baselines integrated by us into the USB codebase.
@@
Pseudo-Labeling Lee et al. (2013) 37.20 17.3 18
Mean Teacher Tarvainen and Valpola (2017) 33.13 16.0 16
$\Pi$-Model Rasmus et al. (2015) 51.13 18.6 19
VAT Miyato et al. (2018) 32.50 13.6 15
MixMatch Berthelot et al. (2019b) 36.31 16.1 17
ReMixMatch Berthelot et al. (2019a) 75.00 20.9 21
AdaMatch Berthelot et al. (2021) 27.59 06.5 05
UDA Xie et al. (2020a) 60.19 19.6 20
FixMatch Sohn et al. (2020) 29.83 10.2 11
FlexMatch Zhang et al. (2022) 28.23 09.4 09
Dash Xu et al. (2021) 30.33 11.7 13
CRMatch Fan et al. (2023) 27.69 07.6 06
CoMatch Li et al. (2021) 27.94 08.1 08
SimMatch Zheng et al. (2022) 27.94 07.6 06
FreeMatch Wang et al. (2023) 28.21 09.7 10
SoftMatch Chen et al. (2023) 27.42 06.4 04
Multihead Co-training Chen et al. (2021) 29.49 12.2 14
MarginMatch Sosea and Caragea (2023) 26.65 02.3 02
SequenceMatch Nguyen (2024) 29.19 11.3 12
CGMatch Cheng et al. (2025) 27.10 03.9 03
MultiMatch Ours 26.37 01.2 01
@@
Appendix H Comparison to Large Language Models
Although a direct comparison between specialized SSL methods and general-purpose LLMs is inherently challenging (e.g., due to potential data contamination and significant disparity in model sizes), we include this analysis to provide a broader context for our proposed approach. These experiments are particularly valuable for assessing the performance of LLMs in specialized classification tasks and for highlighting the continued relevance of specialized, inference-efficient models in a landscape dominated by LLMs. To this end, we evaluate a representative LLM, GPT-4o-mini, on the five NLP datasets used in this work and show the results in Table 8.
Dataset Test Error (%)
IMDB 05.34
AG News 16.26
Amazon Review 36.42
Yahoo! Answer 30.96
Yelp Review 32.96
Table 8: Performance of GPT-4o-mini on the USB datasets.
For the three sentiment classification tasks (IMDB, Amazon Review, and Yelp Review), the performance of GPT-4o-mini is comparable to the Supervised (Full) baseline presented in Table 1, which was trained with large quantities of labeled data (e.g. 250k samples for Amazon and Yelp). This strongly supports the hypothesis that such widely available datasets have been included in LLM training data, making them unsuitable for a fair comparison between SSL and LLMs.
Notably, MultiMatch significantly outperforms GPT-4o-mini on the AG News dataset. To understand the reasons behind this, we perform a quantitative and qualitative analysis of the LLM’s predictions. The confusion matrix shown in Table 9 reveals that the model frequently misclassifies news from the Sci/Tech category as Business, and occasionally World news as Business. A qualitative analysis shows that a large portion of the Sci/Tech news involves corporate activities (e.g., acquisitions, mergers, layoffs, new project announcements). This semantic overlap with the Business category leads to misclassification by the general-purpose LLM. Similarly, some World news regarding inflation rates or trade agreements between governments are also mistakenly classified as Business. These findings highlight the need for specialized training to distinguish between such nuanced categories, where SSL methods like MultiMatch prove effective even under limited labeled data.
True $\backslash$ Pred World Sports Business Sci/Tech
World 0.87 0.03 0.09 0.02
Sports 0.01 0.98 0.01 0.00
Business 0.03 0.00 0.92 0.04
Sci/Tech 0.07 0.01 0.34 0.58
Table 9: Confusion matrix of GPT-4o-mini predictions on AG News.
While these experiments provide valuable insights, we re-emphasize that a direct comparison between our SSL framework and a general-purpose LLM is inherently flawed. First, such a comparison needs to be performed in-depth, using novel datasets not seen during LLM pretraining. Secondly, our framework demonstrates the effectiveness of SSL for training small specialized models using limited labeled data. Such models have a significantly lower inference complexity than LLMs of billions of parameters, making them highly suitable for resource-constrained systems. As detailed in Appendix F, a model trained with SSL preserves the inference complexity of the underlying backbone (e.g., BERT). Therefore, the SSL domain remains highly relevant, providing a complementary approach to LLMs by offering high-performance solutions for specialized, resource-limited tasks.
Appendix I Code and License
The source code for our MultiMatch method and all experimental configurations are publicly available on GitHub at https://github.com/iustinsirbu13/MultiMatch. Our implementation is a public fork of the Unified Semi-supervised Learning (USB) framework Wang et al. (2022) and is released under the MIT license, consistent with the original framework. We have submitted a pull request to the official USB repository to integrate MultiMatch as a new algorithm, thereby contributing our work to the open-source community.